OpenStudy (anonymous):

Consider the equation 3x + 5y = 45. Part 1: On your own paper, graph this equation using the slope-intercept: method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar. (3 points) Part 2: On your own paper, graph this equation using the intercept: method. In the space provided, explain, in words, each step of the procedure you used

OpenStudy (anonymous):

3x+5y=45 =>x/15+y/9=1 first line is written in intercept form. now we get the intercepts as (15,0) and (0,9). now mark these points on graph paper. join these points to get required line. for second question, equation is y+3/x-2=3 y+3=3x-6 3x-y=9 x/3-y/9=1 now we get the intercepts as (3,0) and (0,-9). now mark these points on graph paper. join these points to get required line.

OpenStudy (anonymous):

hope this help :)